13 ) sara co has equity multiplier of 3, what is its debt ratio ( assumme the firm operations are financed by both dept ant equity)
Answer #1

Equity Multiplier = Total Assets / Equity

= 3 / 1

Total Assets = Equity + Liability

3 = 1 + 2

Thus Equity is 1

Liabilty is 2

Debt Equity Ratio = Liability / Equity

= 2 / 1

= 2:1